15. A report indicated that 43% of adults had received a phishing contact (a bogus email that replicates an authentic site
Answer #1

Solution

Given that,

p = 0.43

1 - p = 1 - 0.43 = 0.57

n = 800

\mu\hat p = p = 0.43

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] =  \sqrt [(0.43 * 0.57) / 800 ] = 0.0175

a) P( \hat p < 0.40)

= P[( \hat p - \mu \hat p ) / \sigma \hat p < (0.40 - 0.43) / 0.0175 ]

= P(z < -1.71)

Using z table,

= 0.0436

b) P( 0.37 < \hat p < 0.42 )

= P[(0.37 - 0.43) / 0.0175 < ( \hat p - \mu \hat p ) / \sigma \hat p < (0.42 - 0.43) / 0.0175 ]

= P(-3.43 < z < -0.57)

= P(z < -0.57) - P(z < -3.43)

Using z table,   

= 0.2843 - 0.0003

= 0.2840